OpenStudy (anonymous):

i have to "solve the following system of equations algebraically". but there are two equations and two variables. the first is y=xsquared+2x-1. the second is y=3x+5> I don"t know how to begin.

OpenStudy (anonymous):

the >is supposed to be a period at the end of the second equation.

OpenStudy (anonymous):

set them equal and solve. x^2+2x-1=3x+5 x^2-x-6=0 (x-3)(x+2)=0 x=3, x= -2

OpenStudy (anonymous):

Well, when you are solving systems, you are looking for the points that all the equations have in common. If the points are common, then the y value in one equation will equal the y value in the other equation. Therefore you can set the two expressions for y equal to each other and solve for x.

OpenStudy (anonymous):

still looking at it...what is ^? and what happened to the "y"?

OpenStudy (anonymous):

Ok, you have this: \[y = x^2+2x -1\]\[y=3x+5\] Set those two expressions for y in terms of x equal to each other.

OpenStudy (anonymous):

x^+2x-1=3x+5

OpenStudy (anonymous):

if the two equations are equal, then the 'y' values must be the same. That is why you can set them equal.

OpenStudy (anonymous):

Now from what you have there you can either solve by factoring, or use the quadratic formula to find your x values.

OpenStudy (anonymous):

OK, thank you! i got it!

OpenStudy (anonymous):

no, I don"t have it yet. when I get to the part of the equation of: Xsquared-x-6=0, how do I get rid of the squared X and get to the answer?

OpenStudy (anonymous):

Quadratic equation.. \[ax^2 + bx + c = 0 \implies x = \frac{-(b) \pm \sqrt{(b)^2-4(a)(c)}}{2(a)}\]

OpenStudy (anonymous):

we aren"t doing that yet.....is there another way?

OpenStudy (anonymous):

factor it

OpenStudy (anonymous):

What two numbers add together to make -1 and multiply together to make -6

OpenStudy (anonymous):

(the coefficients of your middle and last terms respectively)

OpenStudy (anonymous):

2 and -3

OpenStudy (anonymous):

Correct. Therefore your equation factors into (x+2)(x-3) = 0

OpenStudy (anonymous):

Which means x can be either -2 (making x+2 = 0) or 3 (making x-3 = 0)

OpenStudy (anonymous):

i don"t see where you got the -1 and 6....i have a -6....at the end of the second one, but a -x on the first

OpenStudy (anonymous):

I tried to apply the same formula on my next equation and got as far as the previous one....I have xsquared-x-2=0. what am I missing?